A thin uniform rod of length l and mass m is swinging freely about a horizontal axis passing through its
end. Its maximum angular speed is ?. Its centre of mass rises to a maximum height of –

Grade:11

2 Answers

Arun
3 years ago
The uniform rod of length l and mass m is swinging about an axis passing through the end. When the centre of mass is raised through h, the increase in potential energy is mgh. This is equal to the kinetic energy = (1/2)Iω2. ⇒ mgh = 1/2(m(l2/3)).ω2 Therefore, h = (l2 . ω2)/6g.
Vikas TU
3 years ago
 
Dear student 
Please refer the clear image , 
The image above does not looks good , it is blurr.
Hope for the best 
Good Luck 
Cheers